OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -3(1.02)^x x g(x) -1 -5 0 -3 1 -1 2 1

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

And here are the answer choices

OpenStudy (anonymous):

The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (jdoe0001):

hmmm how to find the y-intercepts anyway?

OpenStudy (anonymous):

This would help me greatly, How do I find the y intercept on a table

OpenStudy (anonymous):

Yeah, I have the same issue :/

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/finding-intercepts-equation.html <--- what do you think about say f(x)?

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

hmmm did you see how to find the y-intercepts there?

OpenStudy (anonymous):

I'm trying to graph it

OpenStudy (jdoe0001):

well... you don't have to really... but ok

OpenStudy (anonymous):

I could try to find the y intercept on a graph

OpenStudy (anonymous):

Which for f(x) is -3 I think

OpenStudy (anonymous):

For the second line it looks like 1.5

OpenStudy (anonymous):

If it helps, I noticed that the solution of the two lines is 3,3

OpenStudy (jdoe0001):

yeap you could have just set x = 0 btw and \(\bf f(x) = -3(1.02)^{\color{brown}{ x}}\implies f(x) = -3(1.02)^{\color{brown}{ 0}}\implies f(x)=-3\cdot 1\to -3\)

OpenStudy (jdoe0001):

what about the y-intercept of g(x)? any ideas? keep in mind that x = 0 when the graph intercepts the y-axis

OpenStudy (anonymous):

Here is a picture of the graph I made.

OpenStudy (anonymous):

Maybe in some way shape or form it could help

OpenStudy (jdoe0001):

well... let's take a look at your table then \(\large \begin{array}{ccllll} x&g(x) \\\hline\\ -1&-5\\ {\color{brown}{ 0}}&-3\\ 1&-1\\ 2&1 \end{array}\) what do you think the y-intercept would be?

OpenStudy (anonymous):

No clue

OpenStudy (jdoe0001):

well. when x = 0 y = ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

I think

OpenStudy (jdoe0001):

or I should say when x = 0 g(x) = y = ? yeap -3

OpenStudy (jdoe0001):

so now we know that f(x) y-intercept is -3 and g(x)'s y-intercept is -3 so you can infer from that I"d think

OpenStudy (anonymous):

So it would be A, or they are both equal?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

THANK YOU

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

You're a hero :DD

OpenStudy (jdoe0001):

actaully... for that I think I may need a tight outfit with a cape...but I don't have one =)